OpenStudy (anonymous):

Sally needs 50 candy bars to put in her goodie bags for her party. She already has 3 candy bars at home. Each box of candy bars at the store has 8 bars in it. How many boxes of candy bars does she need to buy to reach 50? A. 5 B. 6 C. 7 D. 8

OpenStudy (lucaz):

she needs 50 but already have 3, 50-3=47 each box at the store has 8 bars 47>=8x, solve for x

OpenStudy (anonymous):

5.875 so its a

OpenStudy (lucaz):

sorry, it's 47<=8x

OpenStudy (anonymous):

is it a

OpenStudy (anonymous):

Total candys to be placed are 50 Each box stores 8 Candy bars she had 3 already at home which are capable of storing 8 candies each would be 8 * 3 = 24 So she already stored 24 candys in the 3 boxes which she had in home now left out candies are 50 - 24 = 26 1 box can store 8 Candies ? boxes we need to store 26 Candies = 26 / 8 = 3 (Approx) Total would be 3 in home and 3 she buys from shop which would be 6

OpenStudy (lucaz):

@ganibl she has 3 candy bars, not 3 boxes

OpenStudy (lucaz):

she needs some number of boxes that is greater than or equal to 47

OpenStudy (lucaz):

47 bars

OpenStudy (anonymous):

Got it sorry on that

OpenStudy (lucaz):

1 box = 8 bars 2 boxes = 16 . . 6 boxes = 48 bars

OpenStudy (anonymous):

thank you i know it now

OpenStudy (lucaz):

she needs a minimum of 6 boxes to get 50 bars